# Khatri-Rao Clustering for Data Summarization

### v1

@@ Appendix B Implementation Details
Implementation details of the naïve approach to Khatri-Rao clustering. For the naïve Khatri-Rao clustering baseline used in our experiments, we first run the scikit-learn (Pedregosa et al., 2011) implementation of standard $k$-Means to extract the desired number of cluster centroids. For instance, if the goal is to extract two sets of $h_{1}$ and $h_{2}$ protocentroids, $k$-Means retrieves $h_{1}h_{2}$ clusters. Then, using a coordinate-descent procedure implemented in Python with closed-form updates, we decompose a set of centroids into two smaller sets of protocentroids whose Khatri–Rao product approximates the original centroids. The coordinate-descent procedure alternates between updating the protocentroids of the first and second set, holding the other set fixed. At each step, we use the closed-form updates illustrated in Equation (8) in Section 5. The procedure stops either when a maximum number of iterations is reached ($5000$ by default) or when the total sum of squared differences between the initial centroids and the Khatri-Rao aggregation of the protocentroids becomes smaller than a user-specified threshold ($10^{-4}$ by default). Upon termination of the described coordinate-descent procedure, we have the output sets of protocentroids. The corresponding centroids are readily obtained by aggregating protocentroids. To conclude, we assign each data point to the closest centroid. The implementation is available online in our code repository  [^7] https://github.com/maciap/KhatriRaoClustering/blob/main/scripts/KRkmeansExperimentsLib.py.

### v2

@@ Appendix B Implementation Details
Implementation details of the naïve approach to Khatri-Rao clustering. For the naïve Khatri-Rao clustering baseline used in our experiments, we first run the scikit-learn (scikit-learn) implementation of standard $k$-Means to extract the desired number of cluster centroids. For instance, if the goal is to extract two sets of $h_{1}$ and $h_{2}$ protocentroids, $k$-Means retrieves $h_{1}h_{2}$ clusters. Then, using a coordinate-descent procedure implemented in Python with closed-form updates, we decompose a set of centroids into two smaller sets of protocentroids whose Khatri–Rao product approximates the original centroids. The coordinate-descent procedure alternates between updating the protocentroids of the first and second set, holding the other set fixed. At each step, we use the closed-form updates illustrated in Equation (8) in Section 5. The procedure stops either when a maximum number of iterations is reached ($5000$ by default) or when the total sum of squared differences between the initial centroids and the Khatri-Rao aggregation of the protocentroids becomes smaller than a user-specified threshold ($10^{-4}$ by default). Upon termination of the described coordinate-descent procedure, we have the output sets of protocentroids. The corresponding centroids are readily obtained by aggregating protocentroids. To conclude, we assign each data point to the closest centroid. The implementation is available online in our code repository  [^7] https://github.com/maciap/KhatriRaoClustering/blob/main/scripts/KRkmeansExperimentsLib.py.